OpenStudy (anonymous):

q⁷ x q x q⁻⁹ x q⁻³ A.) q⁻⁶ B.) 1/q⁴ C.) q⁻³ D.) 1/q³

OpenStudy (anonymous):

b

OpenStudy (anonymous):

Thanks.

OpenStudy (anonymous):

welcome again.. ;)

OpenStudy (akashdeepdeb):

Add all the powers because it is multiplied! q^7+1-9-3 q-4 Thus the answer is B) 1/q^4 q^0-4 = q^-4

OpenStudy (uri):

@benetsaidso add all the powers,where its (-) subract it.

OpenStudy (anonymous):

As the basis of the elements is the same (here : q), you can use the form q^x*q^y = q(x+y). This means that you can reduce the problem to an easy summation of powers over q. This comes down to 7+1-9_3 = -4. Remember that a negative exponent translates in 1/exponent, so q^-4 = 1/(q^4), which gives answer b).

OpenStudy (anonymous):

Thanks again guys (: